one integer is 15 more than another. If the product of the integers is -54,find the two integers.
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Try pairs of factors of -54.

Answer by ewatrrr(24785)

x the smaller
x(x+15) = -54
x^2 + 15x + 54 = 0
(x+6)(x+9) = 0
numbers are -6, 9
0r -9, 6